Write an algebraic equation representing the following sentence: The sum of three consecutive odd integers is 45.
Solve the following equation and show your work:
4(x – 12) = -4(14 – x)
Write the following sentence as an algebraic expression: Three less than some number squared.

Answer by stanbon(75887)

Write an algebraic equation representing the following sentence: The sum of three consecutive odd integers is 45.
x + (x+2) + (x+4)
--------------------------
Solve the following equation and show your work:
4(x – 12) = -4(14 – x)
4x-48 = -56 + 4x
-48 = - 56
This is a contradiction.
There is no solution for your equation.
---------------------------
Write the following sentence as an algebraic expression: Three less than some number squared.
x^2-3
----------------
Cheers,
Stan H.